A student has earned scores of 87, 81, and 88 on the first 3 of 4 tests. If the student wants an average(arithmetic mean) of exactly 87, what score must she earn on the fourth test?
A. 85
B. 86
C. 87
D. 92
E. 93

Answer by scott8148(6628)

87+81+88=256 ... 87*4=348 ... 348-256=92